OpenStudy (ja1):

Find the value of:

OpenStudy (ja1):

\[\frac{ 3x }{ \sqrt{x^3} }\]

OpenStudy (anonymous):

what do you want me to do? solve, simplyfie, etc..

OpenStudy (ja1):

I edited the question above :)

OpenStudy (anonymous):

nvrmnd lol i just read it. i will try to do it in a sec

OpenStudy (ja1):

alrighty

OpenStudy (anonymous):

\[\frac{3x \sqrt{x ^{3}}}{ x ^{3} }\]

OpenStudy (ivettef365):

you just have to multiply as follows: 3x \(\sqrt{x^3}\) ----- . ------- \(\sqrt{x^3}\) \(\sqrt{x^3}\)

OpenStudy (anonymous):

so it becomes \[\frac{ 3x ^{2}\sqrt{x} }{ x}\]

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

its over x^3

OpenStudy (ja1):

But where did you get that extra x^3 in the numerator?

OpenStudy (anonymous):

wat u mean u mulitply both sides by square root x^3

OpenStudy (ja1):

ah ok I see what you mean now